OpenStudy (jkbo):

Please solve the blank 8x + = x − 11 Solution: x = -6

OpenStudy (anonymous):

let the blank = y so 8x + y = x - 11 and by applying the value of (X) you can find y -the blank-

OpenStudy (jkbo):

ok so its -6-11? @Hamoody1996

OpenStudy (anonymous):

no y=7x-11 = -(-42) -11 =31 so y = 31

OpenStudy (jkbo):

ok so it would be 31 for the answer? @Hamoody1996

OpenStudy (anonymous):

i think yes